OpenStudy (anonymous):

hi, amm, why cos^2a+cos^2b+ cos^2 c =1 ? i have to explain it in class :/

hero (hero):

I wonder what the a, b, and c represent.

OpenStudy (anonymous):

angles

hero (hero):

I know that, but what is the relationship between angles a, b, and c?

OpenStudy (anonymous):

oh! i really don't know , we are learning about vectors in 3d, and the teacher gave us that homework :/

hero (hero):

Hmm, I found this identity here: http://en.wikibooks.org/wiki/Trigonometry/For_Enthusiasts/Less-Used_Trig_Identities

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

but is isnt true assuming that \(a, b, c\) are the internal angles of a triangle, take for example \(a=b=c=\frac{\pi}{3}\)

hero (hero):

That's why I linked the link @satellite73

OpenStudy (anonymous):

then \(\cos(\frac{\pi}{3})=\frac{1}{2}\) and it is not true that \[3\times\frac{1}{2}=1\]

OpenStudy (anonymous):

ahh so i see

OpenStudy (anonymous):

o.0 im new in this level of math U.U so i don't really know to much about this U.U

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."